### v3

# Levelable graphs

Abstract.
We study a family of positive weighted well-covered graphs, which we call levelable graphs, that are related to a construction of level artinian rings in commutative algebra. A graph $G$ is levelable if there exists a weight function with positive integer values on the vertices of $G$ such that $G$ is well-covered with respect to this weight function. That is, the sum of the weights in any maximal independent set of vertices of $G$ is the same. We describe some of the basic properties of levelable graphs and classify the levelable graphs for some families of graphs, e.g., trees, cubic circulants, Cameron–Walker graphs. We also explain the connection between levelable graphs and a class of level artinian rings. Applying a result of Brown and Nowakowski about weighted well-covered graphs, we show that for most graphs, their edge ideals are not Cohen–Macaulay.
Key words and phrases:
well-covered graphs, level algebra, independence complex
2010 Mathematics Subject Classification:
05C69, 05E40, 13E10
Version: October 24, 2025
1. Introduction
In this paper we introduce levelable graphs, a family of weighted well-covered graphs that can be used to be construct level artinian rings, objects of great interest in commutative algebra. Throughout this paper $G=(V,E)$ denotes a finite simple graph on the vertex set $V=V(G)=\{x_{1},\ldots,x_{n}\}$ and with edge set $E=E(G)$. A subset $W$ of $V$ is an independent set if $e\not\subseteq W$ for all edges $e\in E$. We say $W$ is a maximal independent set if no independent set of $G$ strictly includes $W$. Let ${\rm MaxInd}(G)$ denote the set of maximal independent sets of $G$. In 1970 Plummer [23] called a graph well-covered if every element of ${\rm MaxInd}(G)$ has the same cardinality.
Weighted well-covered graphs, a generalization of well-covered graphs, were introduced in 1998 by Caro, Ellingham, and Ramey [5]. For a fixed field $F$, a weight function on $G$ is a function $w:V(G)\rightarrow F$ that assigns to vertex $x_{i}\in V(G)$ a weight $c_{i}$. Note that we can view any $(c_{1},\ldots,c_{n})\in F^{|V|}$ as a weight function on $G$. A graph $G$ with a weight function $(c_{1},\ldots,c_{n})$ is a weighted well-covered graph if there is a $c\in F$ such that
$\sum_{x_{i}\in W}c_{i}=c~\mbox{for all $W\in{\rm MaxInd}(G)$}.$
In this case, $c$ is called the independence weight of $G$. In this language, a graph is well-covered if $G$ is a weighted well-covered graph with respect to the weight function $(1,\ldots,1)$, and the independence weight is the common cardinality of the maximum independent sets. Associated to any graph $G$ is the vector space ${\rm WCW}(G)$ which consists of all weight functions $(c_{1},\ldots,c_{n})$ that make $G$ a weighted well-covered graph. Understanding the structure of ${\rm WCW}(G)$ has been the focus of one branch of research surrounding weighted well-covered graphs; see, for example, [1, 3, 4, 5, 19, 20, 21, 25].
When the field $F$ is $\mathbb{R}$ or $\mathbb{Q}$, Caro, et al. [5] defined a positive well-covered weighting to be a weight function $(c_{1},\ldots,c_{n})$ that not only makes $G$ weighted well-covered, but $c_{i}>0$ for all $i$. As they remark in [5, p. 653]:
We would guess that determining whether a positive well-covered weighting exists is, in general, a difficult problem.
We are interested in a variation of this “difficult problem” by also requiring the $c_{i}$’s in the weight function to be positive integers. To this end, we make the following definition:
Definition 1.1.
A graph $G$ is a levelable graph if there exists a weight function
$(c_{1},\ldots,c_{n})\in\mathbb{N}^{n}_{>0}$
that makes $G$ a weighted well-covered graph. In this case, we say $G$ is a levelable graph with respect to the weight function $(c_{1},\ldots,c_{n})$.
Our choice of name is inspired by the fact that a levelable graph allows one to construct a level graded artinian ring. Level rings were introduced by Stanley [24] almost 50 years ago as a class of rings that sits “between” Gorenstein rings and Cohen–Macaulay rings. Understanding the properties of level rings has inspired a significant amount of research, including [10, 11, 12, 13, 15, 16, 17, 22], in commutative algebra. By applying a result of Van Tuyl and Zanello [26], we have the following link between graph theory and commutative algebra:
Theorem 1.2.
Let $G=(V,E)$ be a graph, and let $I(G)$ be the edge ideal of $G$ in the ring $R=\mathbb{K}[x_{1},\ldots,x_{n}]$. Then $G$ is a levelable graph with respect to the weight function $(c_{1},\ldots,c_{n})$ if and only if
$R/(I(G)+\langle x_{1}^{c_{1}+1},\ldots,x_{n}^{c_{n}+1}\rangle)$
is a level graded artinian ring.
This connection with combinatorial commutative algebra, as well as the “difficult problem” mentioned above, provides motivation to classify or find families of levelable graphs.
It is immediate that levelable graphs exist – all well-covered graphs are levelable with weight function $(1,\ldots,1).$ The goal of this paper is to identify other families of levelable graphs. Accordingly, we classify all the levelable graphs among the following families:
  1. $\bullet$
    Complete multipartite graphs (see Corollary 2.3),
  2. $\bullet$
    Cameron–Walker graphs (see Theorem 4.2),
  3. $\bullet$
    Co-chordal graphs (see Theorem 5.2),
  4. $\bullet$
    Trees (see Theorem 5.4), and
  5. $\bullet$
    Cubic circulant graphs (see Theorems 6.4 and 6.5).
We also provide constructions of levelable graphs, and obstructions that prevent $G$ from being levelable.
As an interesting by-product of this new connection between weighted well-covered graphs and graded artinian level rings, we can use results about the vector space ${\rm WCW}(G)$ to deduce some results in combinatorial commutative algebra. In particular, Brown and Nowakowski [4] showed that for a random graph (suitably defined), the expected dimension of ${\rm WCW}(G)$ is zero. This implies that for a random graph $G$, there is no choice of positive integers $(a_{1},\ldots,a_{n})$ that makes
$R/(I(G)+\langle x_{1}^{a_{1}},\ldots,x_{n}^{a_{n}}\rangle)$
a level graded artinian ring (see Corollary 7.8). Moreover, Brown and Nowakowski’s result allows us to show that for a random graph $G$, $R/I(G)$ is not Cohen–Macaulay (see Corollary 7.9), complementing results of Docthermann–Newman [8, Corollary 1.4] and Erman-Yang [9, Corollary 7.1].
We have structured this paper so that the first part focuses on the properties of levelable graphs. Only in the last section do we introduce the relevant commutative algebra in order to keep the background on algebra to a minimum. In Section 2 we include some basic properties of levelable graphs and provide a useful lemma that describes an obstruction to a graph being levelable. Section 3 includes some constructions of levelable graphs from a given levelable graph. The sections that follow classify the levelable graphs among some families of graphs. In the final section, we explain the connection to commutative algebra, which motivated our original interest in these graphs, and explain how a result of Brown and Nowakowski has consequences for combinatorial commutative algebra.
2. Basic properties of levelable graphs
We begin by exhibiting some of the basic properties and families of levelable graphs. We first note that we can restrict to connected graphs, that is, for any two vertices of $G$, there is a path in $G$ that connects the two vertices.
Lemma 2.1.
Let $G$ and $H$ be disjoint graphs. Then $G\cup H$ is levelable if and only if $G$ and $H$ are levelable.
Proof.
Let $G$ be a graph on $V(G)=\{x_{1},\ldots,x_{n}\}$ and $H$ a graph on $V(H)=\{y_{1},\ldots,y_{m}\}$. Because the graphs $G$ and $H$ are disjoint, we have
${\rm MaxInd}(G\cup H)=\{W\cup Y~|~W\in{\rm MaxInd}(G),~Y\in{\rm MaxInd}(H)\}.$
(“If”) Suppose that $G$ and $H$ are levelable with respect to the weight functions $(c_{1},\ldots,c_{n})$ and $(d_{1},\ldots,d_{m})$, respectively. Let $c$ be the independence weight of $G$ and $d$ the independence weight of $H$. We claim that $G\cup H$ is a levelable graph with respect to the weight function $(c_{1},\ldots,c_{n},d_{1},\ldots,d_{m})$. Indeed, for any $Z\in{\rm MaxInd}(G\cup H)$, we have $Z=W\cup Y$ with $W\cap Y=\emptyset$ and $W\in{\rm MaxInd}(G)$ and $Y\in{\rm MaxInd}(H)$. Thus
$\sum_{z_{i}\in Z}e_{i}=\sum_{x_{i}\in W}c_{i}+\sum_{y_{j}\in Y}d_{j}=c+d.$
(“Only If”) Suppose that $G\cup H$ is levelable with respect to the weight function
$(e_{1},\ldots,e_{n},e_{n+1},\ldots,e_{n+m})$
with independence weight $e$. We claim $G$ and $H$ are levelable with respect to the weight functions $c=(e_{1},\ldots,e_{n})$ and $d=(e_{n+1},\ldots,e_{n+m})$, respectively. Fix a maximal independent set $Y$ of $H$ and consider any $W_{i}\in{\rm MaxInd}(G)$. Then $W_{i}\cup Y\in{\rm MaxInd}(G\cup H)$ and so
$\sum_{z_{j}\in W_{i}\cup Y}e_{j}=\sum_{z_{j}\in W_{i}}e_{j}+\sum_{z_{j}\in Y}e_{j}=e.$
Rearranging this equation, we have
(2.1) $\sum_{z_{j}\in W_{i}}e_{j}=e-\sum_{z_{j}\in Y}e_{j}.$
Note that since $W_{i}$ is a subset of $\{x_{1},\ldots,x_{n}\}$, the $e_{j}$’s that appear in $\sum_{z_{j}\in W_{i}}e_{i}$ only appear among $\{e_{1},\ldots,e_{n}\}$. Because (2.1) is true for all maximal independent sets of $G$, $(e_{1},\ldots,e_{n})$ is a weight function that makes $G$ a levelable graph with independence weight $e-\sum_{z_{j}\in Y}e_{j}$.
By swapping the roles of $G$ and $H$, the same proof also shows that $H$ is levelable. ∎
As noted in the introduction, all well-covered graphs are levelable. The following lemma allows us to deduce that all complete multipartite graphs, which are not necessarily well-covered, are levelable. Consequently, the set of levelable graphs is strictly larger than the class of well-covered graphs.
Lemma 2.2.
Suppose $G$ is a graph such that the elements of ${\rm MaxInd}(G)$ are pairwise disjoint. Then $G$ is a levelable graph.
Proof.
Suppose ${\rm MaxInd}(G)=\{W_{1},\ldots,W_{t}\}$ with $d_{i}=|W_{i}|$ for $i=1,\ldots,t$. Since every vertex $x_{i}$ appears in some $W_{i}$ and because the $W_{i}$’s are pairwise disjoint, we have that $W_{1}\cup\cdots\cup W_{t}$ is a partition of $V(G)$. Thus, after relabeling, we can assume that
$W_{i}=\{x_{d_{1}+d_{2}+\cdots+d_{i-1}+1},\ldots,x_{d_{1}+d_{2}+\cdots+d_{i}}\}~~\mbox{for $i=1,\ldots,t$},$
where $d_{0}=0.$
Set $d={\rm lcm}(d_{1},\ldots,d_{t})$. A straightforward calculation will now show $G$ is a levelable graph with respect to the weight function
$(\underbrace{d/d_{1},\ldots,d/d_{1}}_{d_{1}},\underbrace{d/d_{2},\ldots,d/d_{2}}_{d_{2}},\ldots,\underbrace{d/d_{t},\ldots,d/d_{t}}_{d_{t}}).$
For integers $d,a_{1},\ldots,a_{d}\geq 1$, the complete $d$-partite graph $K_{a_{1},\ldots,a_{d}}$ is the graph with vertices $V=\{x_{i,j}~|~1\leq i\leq d,~1\leq j\leq a_{i}\}$ and edges of the form $\{x_{i,j},x_{k,l}\}$ for all $1\leq i<k\leq d$. Note that with this notation, the well-known complete graph $K_{d}$ is the graph $K_{1,\ldots,1}$. [^1] This is an abuse of notation, since our notation for a complete multipartite graph would imply that $K_{d}$ should be the $1$-partite graph consisting of $d$ isolated vertices. Moving forward, we will use the standard notation of $K_{d}$ for the complete graph. Since the maximal independent sets of $K_{a_{1},\ldots,a_{d}}$ are the pairwise disjoint sets $W_{i}=\{x_{i,1},\ldots,x_{i,a_{i}}\}$ for $i=1,\ldots,d$, we have the following corollary.
Corollary 2.3.
For all $d,a_{1},\ldots,a_{d}\geq 1$, the complete $d$-partite graph $K_{a_{1},\ldots,a_{d}}$ is levelable.
The independence number of $G$, denoted $\alpha(G)$, is the size of the largest maximal independent set of $G$. If $\alpha(G)$ is small, then $G$ must also be levelable.
Theorem 2.4.
If $G$ is a graph on $\{x_{1},\dots,x_{n}\}$ with $\alpha(G)\leq 2$, then $G$ is levelable.
Proof.
If $\alpha(G)=1$, then $G=K_{a}$ is the complete $a$-partite graph (or simply, complete graph) and is levelable by Theorem 2.3.
Suppose that $\alpha(G)=2$. If all $W\in{\rm MaxInd}(G)$ have $|W|=2$, then $G$ is well-covered, and thus levelable. So, suppose that $|W|=1$ for at least one $W\in{\rm MaxInd}(G)$. Let ${\rm MaxInd}(G)=\{W_{1},\ldots,W_{t}\}$. After relabeling we can assume that $|W_{1}|=\cdots=|W_{s}|=1$ and $|W_{s+1}|=\cdots=|W_{t}|=2$. Note that if $|W_{j}|=1$, then $W_{j}\cap W_{k}=\emptyset$ for all $j\neq k$. So, we can relabel again so that $W_{i}=\{x_{i}\}$ for $i=1,\ldots,s$, and $W_{s+1}\cup\cdots\cup W_{t}=\{x_{s+1},\ldots,x_{n}\}$. Then $G$ is levelable with respect to the weight function
$c=(\underbrace{2,\ldots,2}_{s},\underbrace{1,\ldots,1}_{n-s})$
since $\sum_{x_{i}\in W_{i}}c_{i}=2$ if $|W_{i}|=1$ and $\sum_{x_{i}\in W_{i}}c_{i}=2$ if $|W_{i}|=2$. ∎
We let $C_{n}$ denote the $n$-cycle on $n$-vertices with $n\geq 3$. For any graph $G$, we write $G^{c}$ for the graph complement of $G$, that is, the graph with the same vertex set as $G$, but edge set $\{\{x_{i},x_{j}\}:\{x_{i},x_{j}\}\not\in E(G)\}$. When $n=3$, $C_{n}^{c}$ is three disjoint vertices, and thus clearly levelable. On the other hand, $\alpha(C_{n}^{c})=2$ if $n\geq 4$. Consequently, we have the following corollary:
Corollary 2.5.
For any $n\geq 3$, $C_{n}^{c}$ is levelable.
The next example exhibits a non-levelable graph with $\alpha(G)=3$. In fact, this example illustrates that levelable graphs are not preserved by taking induced subgraphs.
Example 2.6.
Let $P_{n}$ denote the path graph of length n1n-1 on $n\geq 2$ vertices, that is, the graph with vertex set $\{x_{1},\ldots,x_{n}\}$ and edge set $\{\{x_{i},x_{i+1}\}~:~1\leq i\leq n-1\}$. We show that $P_{5}$ (see Figure 1) is not levelable. We see that ${\rm MaxInd}(P_{5})=\{\{x_{1},x_{3},x_{5}\},\{x_{1},x_{4}\},\{x_{2},x_{4}\},\{x_{2},x_{5}\}\}$, and thus $\alpha(G)=3$.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$
Figure 1. The path $P_{5}$ which is non-levelable
Suppose that a weight function $(c_{1},\ldots,c_{5})$ makes $P_{5}$ levelable. Then
$(c_{1}+c_{3}+c_{5})+(c_{2}+c_{4})=(c_{1}+c_{4})+(c_{2}+c_{5}).$
Consequently $c_{3}=0$, contradicting the fact that the entries of $(c_{1},\ldots,c_{5})$ are all positive.
The graph given in Figure 2 is a well-covered graph that has $P_{5}$ as an induced subgraph.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$y_{1}$$y_{2}$$y_{3}$$y_{4}$$y_{5}$
Figure 2. A levelable graph with $P_{5}$ as an induced subgraph
Consequently, the levelable property is not preserved by taking induced subgraphs.
Embedded in Example 2.6 is a condition on the elements of ${\rm MaxInd}(G)$ that prevents $G$ from being levelable. We make this condition explicit in the next lemma.
Lemma 2.7.
Suppose there exist $F_{1},F_{2},F_{3},F_{4}\in{\rm MaxInd}(G)$ such that $F_{3}\cup F_{4}\subsetneq F_{1}\cup F_{2}$ and $F_{3}\cap F_{4}=\emptyset$. Then G is not levelable.
Proof.
Suppose towards a contradiction that ($c_{1},\dots,c_{n}$) is a weight function for $G$ with independence weight $c$. Then $\sum_{x_{j}\in F_{1}}c_{j}+\sum_{x_{j}\in F_{2}}c_{j}=\sum_{x_{j}\in F_{3}}c_{j}+\sum_{x_{j}\in F_{4}}c_{j}=2c$. Since $c_{j}>0$ for all $j$, and $F_{3}\cup F_{4}\subsetneq F_{1}\cup F_{2}$, we must have
$\sum_{x_{j}\in F_{3}\cup F_{4}}c_{j}<\sum_{x_{j}\in F_{1}\cup F_{2}}c_{j}\leq\sum_{x_{j}\in F_{1}}c_{j}+\sum_{x_{j}\in F_{2}}c_{j}=2c.$
But then since $F_{3}\cap F_{4}=\emptyset$, we have
$2c=\sum_{x_{j}\in F_{3}}c_{j}+\sum_{x_{j}\in F_{4}}c_{j}=\sum_{x_{j}\in F_{3}\cup F_{4}}c_{j}<2c,$
which is impossible. ∎
This next result, which can also be deduced from Theorem 5.4 given later in the paper, demonstrates how to apply Lemma 2.7.
Corollary 2.8.
The path graph $P_{n}$ is levelable if and only if $n\in\{2,3,4\}$.
Proof.
When $n=2$ or $4$, the graph $P_{n}$ is well-covered, and thus levelable. When $n=3$, the weight function $(1,2,1)$ makes $P_{3}$ levelable.
Now suppose that $n\geq 5$. Suppose that $n=2m+1$ is odd. Then
$F_{1}$ $=$ $\{x_{1},x_{3},\ldots,x_{2m-1},x_{2m+1}\},~~F_{2}=\{x_{2},x_{4},\ldots,x_{2m}\},$
$F_{3}$ $=$ $\{x_{1},x_{4},x_{6},\ldots,x_{2m}\}~~\mbox{and}~~F_{4}=\{x_{2},x_{5},x_{7},\ldots,x_{2m+1}\}$
are all elements of ${\rm MaxInd}(G)$. The result now follows from Lemma 2.7.
Similarly, if $n=2m$ is even, we have $m\geq 3$, and we consider the sets:
$F_{1}$ $=$ $\{x_{1},x_{3},\ldots,x_{2m-1}\},F_{2}=\{x_{2},x_{4},\ldots,x_{2m}\},$
$F_{3}$ $=$ $\{x_{1},x_{4},x_{6},\ldots,x_{2m}\}~~\mbox{and}~~F_{4}=\{x_{2},x_{5},x_{7},\ldots,x_{2m-1}\}$
and again use Lemma 2.7. ∎
Remark 2.9.
While our paper has focused on independent sets, we also could have of used the notion of vertex covers to define levelable graphs. A subset $W\subseteq V(G)$ is called a vertex cover if $e\cap W\neq\emptyset$ for all $e\in E(G)$. A vertex cover is the complement of an independent set, that is, $W$ is a vertex cover of $G$ if and only if $V\setminus W$ is an independent set. Maximal independent sets correspond to minimal vertex covers.
If we let ${\rm MinVC}(G)$ be the collection of minimal vertex covers, then the following definition for levelable graphs is equivalent to Definition 1.1: a graph $G$ is levelable if there exists a weight function $(d_{1},\ldots,d_{n})\in\mathbb{N}^{n}_{>0}$ and some constant $d$ such that
$\sum_{x_{i}\in W}d_{i}=d~~\mbox{for all $W\in{\rm MinVC}(G)$}.$
To see that these definitions are equivalent, set $D=d_{1}+\cdots+d_{n}$. Then
$\sum_{x_{i}\in V\setminus W}d_{i}=D-\sum_{x_{i}\in W}d_{i}=D-d.$
But as $W$ runs through all elements of ${\rm MinVC}(G)$, $V\setminus W$ runs through all elements of ${\rm MaxInd}(G)$. So $G$ is levelable with respect to Definition 1.1. By simply reversing this argument, we can show Definition 1.1 is equivalent to this reformulation.
3. Constructions
In this section, we introduce basic techniques to construct a new levelable graph from a levelable graph.
Definition 3.1.
Let $G$ be a finite graph and fix a vertex $x$ of $G$.
  • The graph $G^{x}$, called the duplication of $G$ at $x$, is the finite graph obtained by adding a new vertex $y$ to $G$ with
    $E(G^{x})=E(G)\cup\{\{y,b\}:\{x,b\}\in E(G)\}.$
  • The graph $G^{[x]}$, called the expansion of $G$ at $x$, is the finite graph obtained by adding a new vertex $y$ to $G$ with
    $E(G^{[x]})=E(G^{x})\cup\{\{x,y\}\}.$
Theorem 3.2.
Suppose that $G$ is a levelable graph and fix a vertex $x$ of $G$. Then both $G^{x}$ and $G^{[x]}$ are levelable.
Proof.
Let $G$ be a levelable graph on $V=\{x_{1},\ldots,x_{n}\}$ with respect to a weight function $(c_{1},\ldots,c_{n})$. Set $x=x_{n}$ and $y=x_{n+1}$.
Let $W$ be a maximal independent set of $G^{x}$. If $x\not\in W$, then $W$ is a maximal independent set of $G$. If $x\in W$, then $y\in W$ and $W\setminus\{y\}$ is a maximal independent set of $G$. Hence $G^{x}$ is levelable with respect to the weight function
$(2c_{1},\cdots,2c_{n-1},c_{n},c_{n}).$
A maximal independent set $W$ of $G^{[x]}$ is either a maximal independent set of $G$ or $(W\setminus\{x\})\cup\{y\}$, where $W$ is a maximal independent set of $G$ with $x\in W$. Hence $G^{[x]}$ is levelable with respect to the weight function
$(c_{1},\cdots,c_{n-1},c_{n},c_{n}).$
Example 3.3.
The pentagon is levelable since it is a well-covered graph. Then repeated applications of the duplication technique shows that the graph of Figure 3 is levelable.
Figure 3. A pentagon with duplications
Corollary 3.4.
Let $G$ be a levelable graph on $\{x_{1},\ldots,x_{n}\}$ with respect to a weight function $(c_{1},c_{2},\ldots,c_{n})$. Then for any positive integers $r_{1},\ldots,r_{n}$, there is a graph $G^{\prime}$ which is levelable with respect to the weight function
(3.1) $(\underbrace{c_{1},c_{1},\ldots,c_{1}}_{r_{1}},\underbrace{c_{2},c_{2},\ldots,c_{2}}_{r_{2}},\ldots,\underbrace{c_{n},c_{n},\ldots,c_{n}}_{r_{n}}).$
Proof.
Let $r_{i}$ be the number of $c_{i}$ in (3.1). Then doing $r_{i}-1$ expansions of each vertex $x_{i}$ yields the desired graph $G^{\prime}$. ∎
Remark 3.5.
Let $G$ be a graph consisting of $n$ isolated vertices. Since $\{x_{1},\ldots,x_{n}\}$ is the unique maximal independent set of $G$, it follows that $G$ is levelable with respect to any weight function $(c_{1},\ldots,c_{n})$. In particular, given any sequence $(c_{1},\ldots,c_{n})$ of positive integers, there is a levelable graph on $\{x_{1},\ldots,x_{n}\}$ with respect to $(c_{1},\ldots,c_{n})$. Let $n=3$. Then the connected graphs are the triangle and the path of length $2$. Hence given a sequence $(c_{1},c_{2},c_{3})$ of positive integers with $c_{1}\leq c_{2}\leq c_{3}$, there is a connected levelable graph on $\{x_{1},x_{2},x_{3}\}$ with respect to $(c_{1},c_{2},c_{3})$ if and only if either $c_{1}=c_{2}=c_{3}$ or $c_{3}=c_{1}+c_{2}$. It would, of course, be of interest to classify all possible weight functions for connected levelable graphs. Note that if $G$ is a levelable graph on $\{x_{1},\ldots,x_{n}\}$ with respect to a weight function $(c_{1},\ldots,c_{n})$, then $G$ is also levelable with respect to the weight function $(mc_{1},\ldots,mc_{n})$, where $m$ is a positive integer.
Definition 3.6.
Let $G$ be a finite connected graph on $V(G)=\{x_{1},\ldots,x_{n}\}$. Given finite graphs $H_{1},\ldots,H_{n}$ with each $V(H_{i})=\{z^{(i)}_{1},\ldots,z^{(i)}_{r_{i}}\}$, where $V(G),V(H_{1}),\ldots,V(H_{n})$ are pairwise disjoint, we introduce the finite graph $G(H_{1},\ldots,H_{n})$ on $V(G)\cup V(H_{1})\cup\cdots\cup V(H_{n})$ whose set of edges is
$E(G)\cup E(H_{1})\cup\cdots\cup E(H_{n})\cup\left(\,\bigcup_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq r_{i}\end{subarray}}\{\{x_{i},z^{(i)}_{j}\}\}\right)$.
Theorem 3.7.
With the notation as in Definition 3.6, the graph $G(H_{1},\ldots,H_{n})$ is levelable if and only if each $H_{i}$ is levelable.
In particular, if each $H_{i}$ is a complete graph $K_{r_{i}}$, then $G(H_{1},\ldots,H_{n})$ is levelable.
Proof.
Let $\Gamma=G(H_{1},\ldots,H_{n})$. A maximal independent set of $\Gamma$ consists of an independent set $W_{0}$ of $G$ together with maximal independent sets $W_{i}$ of $H_{i}$ for each $i$ such that $x_{i}\not\in W_{0}$.
(“If”) Suppose that each $H_{i}$ is levelable with respect to a weight function
$(c(z^{(i)}_{1}),\ldots,c(z^{(i)}_{r_{i}}))$.
Let $d_{i}=\sum_{z^{(i)}_{j}\in W}c(z^{(i)}_{j})$ be its independence weight, where $W$ is a maximal independent set of $H_{i}$. Let $c(x_{i})=d_{i}$ for $1\leq i\leq n$. It can be checked that $\Gamma$ is levelable with respect to the weight function
(3.2) $(c(x_{1}),\ldots,c(x_{n}),c(z_{1}^{(1)}),\ldots,c(z_{r_{n}}^{(n)})).$
(“Only If”) Suppose that $\Gamma$ is levelable with respect to a weight function of the form (3.2). Let $W$ be a maximal independent set of $\Gamma$ with $x_{i}\not\in W$ and $W_{i}$ a maximal independent set of $H_{i}$ with $W_{i}\subseteq W$. Let $W^{\prime}_{i}$ be a maximal independent set of $H_{i}$ with $W_{i}\neq W_{i}^{\prime}$. Since $(W\setminus W_{i})\cup W^{\prime}_{i}$ is a maximal independent set of $\Gamma$, it follows that $H_{i}$ is levelable with respect to the weight function $(c(z^{(i)}_{1}),\ldots,c(z^{(i)}_{r_{i}}))$. ∎
Corollary 3.8.
Given positive integers $c_{1},\ldots,c_{n}$ and positive integers $r_{1},\ldots,r_{n}$ with $r_{i}\geq 2$ for all $i$, there is a connected levelable graph on $r_{1}+\cdots+r_{n}$ vertices with respect to the weight function
$(\underbrace{c_{1},\ldots,c_{1}}_{r_{1}},\underbrace{c_{2},\ldots,c_{2}}_{r_{2}},\ldots,\underbrace{c_{n},\ldots,c_{n}}_{r_{n}}).$
Proof.
Let $P_{n}$ be the path on $\{x_{1},\ldots,x_{n}\}$ and let $K_{r_{i}-1}$ be the complete graph on the vertex set $\{z_{1}^{(i)},\ldots,z_{r_{i}-1}^{(i)}\}$. By Theorem 3.7, the graph $G=P_{n}(K_{r_{1}-1},\ldots,K_{r_{n}-1})$ is levelable. A maximal independent set of $G$ is of the form $A\cup B$, where $A$ is an independent set of $P_{n}$, $B$ is a subset of $\bigcup_{i=1}^{n}\{z_{1}^{(i)},\ldots,z_{r_{i}-1}^{(i)}\}$, and for each $x_{i}\not\in A$, $|B\cap\{z_{1}^{(i)},\ldots,z_{r_{i}-1}^{(i)}\}|=1$, that is, $B$ contains exactly one of the vertices of $K_{r_{i}-1}$. In particular, $G$ is well-covered. Because each maximal independent set of $G$ contains exactly one vertex of $\{x_{i},z_{1}^{(i)},\ldots,z_{r_{i}-1}^{(i)}\}$ for each $i$, the graph $G$ is then levelable with respect to the weight function
$(c(x_{1}),c(z_{1}^{(1)}),\ldots,c(z_{r_{1}-1}^{(1)}),c(x_{2}),c(z_{1}^{(2)}),\ldots,c(z_{r_{2}-1}^{(2)}),\ldots,c(x_{n}),c(z_{1}^{(n)}),\ldots,c(z_{r_{n}-1}^{(n)}))\\ =(\underbrace{c_{1},\ldots,c_{1}}_{r_{1}},\underbrace{c_{2},\ldots,c_{2}}_{r_{2}},\ldots,\underbrace{c_{n},\ldots,c_{n}}_{r_{n}}).$
Remark 3.9.
In the proof of Corollary 3.8, since $G$ is well-covered, it follows that $G$ is also levelable with respect to the weight function $(1,\ldots,1)$.
Corollary 3.10.
Given positive integers $c_{1},\ldots,c_{n}$, there is a connected levelable graph on $V=\{x_{1},\ldots,x_{n},\ldots,x_{N}\}$ with respect to the weight function
$(c_{1},\ldots,c_{n},1,\ldots,1)\in{\mathbb{Z}}^{N},$
where
$N=n+\sum_{i=1}^{n}c_{i}.$
Proof.
Let $P_{n}$ be the path on $\{x_{1},\ldots,x_{n}\}$. We introduce the graph $G$ on
$\{x_{1},\ldots,x_{n}\}\cup\{y_{j}^{(i)}:1\leq i\leq n,1\leq j\leq c_{i}\}$
which is obtained by adding edges $\{x_{i},y_{1}^{(i)}\},\ldots,\{x_{i},y_{c_{i}}^{(i)}\}$ to $P_{n}$. Theorem 3.7 says that $G$ is levelable. A maximal independent set of $G$ is of the form $A\cup B$, where $A$ is an independent set of $P_{n}$ and $B=\{y_{j}^{(i)}:x_{i}\not\in A,1\leq j\leq c_{i}\}$. Hence $G$ is levelable with respect to the weight function
$(c(x_{1}),\ldots,c(x_{n}),c(y^{(1)}_{1}),\ldots,c(y^{(n)}_{c_{n}}))=(c_{1},\ldots,c_{n},1,1,\ldots,1)\in{\mathbb{Z}}^{N}.$
4. Cameron–Walker graphs
We now turn to the discussion of the problem of which Cameron–Walker graphs are levelable. A Cameron–Walker graph [6, 18] is a finite graph consisting of a connected bipartite graph with vertex partition $U\cup V$ such that there is at least one leaf (or pendant edge) attached to each $x\in U$ and that there may be possibly some pendant triangles attached to each $y\in V$. A vertex $y\in V$ is called exceptional if there is no pendant triangle attached to $y$. It will be assumed that, for each exceptional $y\in V$, there exists $x,x^{\prime}\in U$ with $x\neq x^{\prime}$ for which $\{x,y\},\{x^{\prime},y\}\in E(G)$.
Example 4.1.
The Cameron–Walker graph $G$ of Figure 4 is not levelable. To see why this is true, suppose that $G$ is levelable with respect to a weight function $(c_{1},\ldots,c_{10})$. Since $\{x_{1},x_{2},x_{3},x_{7},x_{8}\}$ and $\{x_{1},x_{2},x_{6},x_{7}\}$ are maximal independent sets of $G$, one has $c_{3}+c_{8}=c_{6}$. Furthermore, since $\{x_{3},x_{4},x_{5},x_{9}\}$ and $\{x_{4},x_{5},x_{6},x_{9}\}$ are maximal independent sets of $G$, one has $c_{3}=c_{6}$. Hence $c_{8}=0$, a contradiction.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$x_{8}$$x_{9}$$x_{10}$
Figure 4. A non-levelable Cameron–Walker graph
We now come to a classification of levelable Cameron–Walker graphs.
Theorem 4.2.
A Cameron–Walker graph is levalable if and only if it has no exceptional vertex.
Proof.
First, the “If” part follows from Theorem 3.7. To prove the “Only If” part, suppose that $G$ is a Cameron–Walker graph with an exceptional vertex. Let $U=\{x_{1},\ldots,x_{a}\}$ and $V=\{y_{1},\ldots,y_{b}\}$. Let $A_{i}$ denote the set of vertices $z$ of $G$ for which $\{x_{i},z\}$ is a leaf attached to $x_{i}$ for $1\leq i\leq a$ and $B_{j}$ the set of edges $\{v,w\}$ for which $G|_{\{y_{j},v,w\}}$ is a pendant triangle attached to $y_{j}$ for $1\leq j\leq b$. Each $A_{i}$ is nonempty and some $B_{j}$ might be empty. Let $A_{i}=\{z_{1}^{(i)},\ldots,z_{q_{i}}^{(i)}\}$ and $B_{j}=\{\{v_{1}^{(j)},w_{1}^{(j)}\},\ldots,\{v_{r_{j}}^{(j)},w_{r_{j}}^{(j)}\}\}$. Let $V^{\prime}$ denote the set of exceptional vertices. Without loss of generality, let $y_{1}$ be an exceptional vertex and $\{x_{1},y_{1}\}\in E(G)$. Now,
$(A_{1}\cup\cdots\cup A_{a})\cup V^{\prime}\cup\left(\,\bigcup_{y_{j}\not\in V^{\prime}}\{v_{1}^{(j)},\ldots,v_{r_{j}}^{(j)}\}\right)$
is a maximal independent set of $G$. Furthermore,
$\{x_{1}\}\cup(A_{2}\cup\cdots\cup A_{a})\cup\{y_{j}\in V^{\prime}:\{x_{1},y_{j}\}\not\in E(G)\}\cup\left(\,\bigcup_{y_{j}\not\in V^{\prime}}\{v_{1}^{(j)},\ldots,v_{r_{j}}^{(j)}\}\right)$
[…]
(4.1) $c(z^{(1)}_{1})+\cdots+c(z^{(1)}_{q_{1}})+\sum_{y_{j}\in V^{\prime},\,\{x_{1},y_{j}\}\in E(G)}c(y_{j})=c(x_{1}).$
On the other hand,
$U\cup\left(\,\bigcup_{y_{j}\not\in V^{\prime}}\{v_{1}^{(j)},\ldots,v_{r_{j}}^{(j)}\}\right)$
is a maximal independent set $G$. Furthermore, by using the assumption that, for each $y\in V^{\prime}$, there exists $x,x^{\prime}\in U$ with $x\neq x^{\prime}$ for which $\{x,y\},\{x^{\prime},y\}\in E(G)$, it follows that
$A_{1}\cup\{x_{2},\ldots,x_{a}\}\cup\left(\,\bigcup_{y_{j}\not\in V^{\prime}}\{v_{1}^{(j)},\ldots,v_{r_{j}}^{(j)}\}\right)$
is a maximal independent set of $G$. Hence
(4.2) $c(x_{1})=c(z^{(1)}_{1})+\cdots+c(z^{(1)}_{q_{1}})$.
Finally, by (4.1) and (4.2), one has
$\sum_{y_{j}\in V^{\prime},\,\{x_{1},y_{j}\}\in E(G)}c(y_{j})=0,$
a contradiction. ∎
Suppose that $G$ is a Cameron–Walker graph with no exceptional vertex and work in the same notation as in the proof of Theorem 4.2. Then $G$ is levelable with respect to the weight function
$(c(x_{1}),\ldots,c(y_{b}),c(z^{(1)}_{1}),\ldots)=(q_{1},\ldots,q_{a},r_{1},\ldots r_{b},1,1,\ldots,1,1).$
5. Chordal graphs
Let $\Delta$ be a simplicial complex (the formal definitions are postponed until Section 7). A vertex $x$ of $\Delta$ is called free if $x$ belongs to exactly one facet. A facet $F$ of $\Delta$ is said to be a leaf of $\Delta$ if either $\Delta=\langle F\rangle$, or there exists a facet $G$ of $\Delta$ with $G\neq F$, called a branch of $F$, for which $H\cap F\subseteq G\cap F$ for all facets $H$ of $\Delta$ with $H\neq F$. Each vertex belonging to $F\setminus G$ is free. A labeling $F_{1},\ldots,F_{s}$ of the facets of $\Delta$ is called a leaf order if, for each $1<i\leq s$, the facet $F_{i}$ is a leaf of the subcomplex $\langle F_{1},\ldots,F_{i}\rangle$. A simplicial complex possessing a leaf order is called a quasi-forest. We refer the reader to [14, Chapter 9] for the detailed information.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$F_{1}$$F_{2}$$F_{3}$$F_{4}$
Figure 5. A leaf order
Lemma 5.1.
Let $\Delta$ be a quasi-forest on $\{x_{1},\ldots,x_{n}\}$ and suppose that $F_{1},\ldots,F_{s}$ is a leaf order. Then there is a sequence $(c_{1},\ldots,c_{n})$ of positive integers and an integer $c$ for which $\sum_{i\in F_{j}}c_{i}=c$ for all $F_{j}$.
Proof.
Let $x_{t},x_{t+1},\ldots,x_{n}$ be the free vertices of $\Delta$ in $F_{s}$. Let $G=F_{k}$ be a branch of the leaf $F_{s}\in\Delta$. The simplicial complex $\langle F_{1},\ldots,F_{s-1}\rangle$ on $\{x_{1},\ldots,x_{t-1}\}$ is a quasi-forest with a leaf order $F_{1},\ldots,F_{s-1}$. Thus, by using induction on $s$, there is a sequence $(c_{1},\ldots,c_{t-1})$ of positive integers and an integer $c$ for which $\sum_{x_{i}\in F_{j}}c_{i}=c$ for all $F_{j}$ with $1\leq j<s$. Let $\sum_{x_{i}\in(G\cap F_{s})}c_{i}=c^{\prime}$. One may assume that $c-c^{\prime}>n-t+1$ since we can replace $(c_{1},\ldots,c_{t-1})$ and $c$ with $(dc_{1},\ldots,dc_{t-1})$ and $dc$ for any positive integer $d\geq 1$. Let $c_{t},\ldots,c_{n}$ be positive integers with $c_{t}+\cdots+c_{n}=c-c^{\prime}$. Note that $F_{s}\setminus G=\{x_{t},x_{t+1},\ldots,x_{n}\}$. It then follows that $(c_{1},\ldots,c_{n})$ has the required property. ∎
Recall from [14, Theorem 9.2.12] that the clique complex of a finite graph $G$ is a quasi-forest if and only if $G$ is a chordal graph. A co-chordal graph is a finite graph for which the complement graph $G^{c}$ of $G$ is chordal.
Theorem 5.2.
A co-chordal graph is levelable.
Proof.
Let $G$ be a co-chordal graph on $\{x_{1},\ldots,x_{n}\}$. Then the set of maximal independent sets of $G$ is the set of maximal cliques of the chordal graph $G^{c}$. Thus the desired result follows from Lemma 5.1. ∎
Example 5.3.
The chordal graph $G$ of Figure 5 is levelable, since $G$ is co-chordal. It follows from Theorem 4.2 that the path of $P_{5}$ of Figure 6 is non-levelable, thus giving a proof different from Example 2.6. On the other hand, by Theorem 3.7, the tree of Figure 7 is levelable.
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$
Figure 6. A non-levelable tree
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$
Figure 7. A levelable tree
It is a reasonable project to classify all levelable chordal graphs. As a first step in achieving this goal, one must classify all levelable trees. We now turn our attention to this goal.
Recall that a free vertex of a tree is a vertex which belongs to exactly one edge. Every tree has at least two free vertices. The distance of vertices $x$ and $y$ in a tree is the length of the unique path between $x$ and $y$ in the tree. Note that in a tree a path between $x$ and $y$ is unique.
Theorem 5.4.
A tree is levelable if and only if each non-free vertex is adjacent to at least one free vertex.
Proof.
Theorem 3.7 guarantees that “If” is true. To show that “Only If” is true, suppose that a tree $G$ is levelable and that there is a non-free vertex $x$ which is adjacent to no free vertex. Let $G_{1},\ldots,G_{q}$ with $q\geq 2$ denote the connected components of $G\setminus\{x\}$. Each $G_{i}$ has at least one edge.
Let $q>2$. Taking a maximal independent set $W$ of $G_{q}$ which does not contain the vertex of $G_{q}$ which is adjacent to $x$, it follows that $G\setminus G_{q}$ is levelable. Since $q>2$, the vertex $x$ cannot be a free vertex of $G\setminus G_{q}$. Thus working with induction on the number of vertices of a tree guarantees the existence of a free vertex of $G\setminus G_{q}$ which is adjacent to $x$, a contradiction. Hence $q=2$.
Let $q=2$. Let $A$ denote the set of vertices $y$ of $G$ for which the distance of $y$ and $x$ in $G$ is $2$. Thus both $A\cap G_{1}$ and $A\cap G_{2}$ are nonempty. Let $y_{i}$ denote the vertex of $G_{i}$ which is adjacent to $x$ for $i=1,2$. Let $W$ be a maximal independent set of $G_{1}$ which contains $A\cap G_{1}$ and $W^{\prime}$ a maximal independent set of $G_{2}$ which contains $A\cap G_{2}$. Let $U$ be a maximal independent set of $G_{1}$ which contains $y_{1}$ and $U^{\prime}$ a maximal independent set of $G_{2}$ which contains $y_{2}$. Then
$W\cup W^{\prime}\cup\{x\},\,\,\,U\cup U^{\prime},\,\,\,W^{\prime}\cup U,\,\,\,W\cup U^{\prime}$
are maximal independent sets of $G$. Let $\{x_{1},\ldots,x_{n}\}$ be the vertex set of $G$ and suppose that $G$ is levelable with respect to a weight function $(c_{1},\ldots,c_{n})$. Let $a=\sum_{x_{i}\in W}c_{i}$, $b=\sum_{x_{i}\in W^{\prime}}c_{i}$, $c=\sum_{x_{i}\in U}c_{i}$ and $d=\sum_{x_{i}\in U^{\prime}}c_{i}$. Let $x=x_{1}$. It then follows that the positive integers
$c_{1}+a+b,\,\,\,c+d,\,\,\,b+c,\,\,\,a+d$
are equal. Hence $a=c$. Thus $c_{1}=0$, a contradiction. ∎
Remark 5.5.
Levit and Tankus [20] described a polynomial time algorithm to determine if a graph that is $C_{4},C_{5}$ and $C_{6}$-free is a weighted well-covered graph. Since trees do not have any cycles, it can be shown that the algorithm of [20] produces the previous result, namely, that a tree is levelable if and only if every non-free vertex is adjacent to at least on free vertex.
Definition 5.6.
A graph $G$ is called a caterpillar graph if $G$ is a tree and if all the vertices of $G$ are within distance one of a central path. Each vertex that is not on the central path is called a leg of $G$.
Figure 8. A caterpillar graph
One will assume that each end vertex of a central path has at least one leg. If not, then it can be regarded as a leg of a shorter central path.
Corollary 5.7.
Let $G$ be a caterpillar graph. Then $G$ is levelable if and only if every vertex of a central path has at least one leg.
Theorem 5.4 gives an alternative proof to Corollary 2.8.
Corollary 5.8.
The path $P_{n}$ is levelable if and only if $n\in\{2,3,4\}$.
Definition 5.9.
Let $q\geq 3$ be an integer and $n_{1},\ldots,n_{q}$ positive integers. Let $P_{i}$ denote the path of length $n_{i}$ on the vertex set $V_{i}=\{x,x_{1}^{(i)},\ldots,x_{n_{i}}^{(i)}\}$ for $1\leq i\leq q$. We assume $V_{i}\cap V_{j}=\{x\}$ for $i\neq j$. Let $G(n_{1},\ldots,n_{q})=P_{1}\cup\cdots\cup P_{q}$. We call $G(n_{1},\ldots,n_{q})$ a big star graph with a center $x$.
Figure 9. The big star graph $G(1,2,2,3)$
Corollary 5.10.
A big star graph $G(n_{1},\ldots,n_{q})$ is levelable if and only if each $n_{i}\leq 2$ and there is at least one $j$ with $n_{j}=1$.
6. Circulant graphs
In this section, we concern ourselves with classifying circulant levelable graphs.
Definition 6.1.
Let $n\geq 2$ and $S\subseteq\{1,\dots,\lfloor\frac{n}{2}\rfloor\}$. The circulant graph $C_{n}(S)$ is the graph on the vertex set $\{x_{1},\dots,x_{n}\}$ such that $\{x_{i},x_{j}\}\in E(C_{n}(S))$ precisely when $|i-j|\in S$ or $n-|i-j|\in S$.
Figure 10 shows an example of the circulant graph $C_{10}(2,5)$. Theorem 2.3 and Corollary 2.5 already classify when the circulant graphs $C_{n}(1,2,\ldots,\lfloor\frac{n}{2}\rfloor)$ and $C_{n}(2,\ldots,\lfloor\frac{n}{2}\rfloor)$ are levelable since $C_{n}(1,2,\ldots,\lfloor\frac{n}{2}\rfloor)\cong K_{n}$ and $C_{n}(2,\ldots,\lfloor\frac{n}{2}\rfloor)\cong C_{n}^{c}$. We now classify two further families of circulant graphs.
We first consider the case when $S=\{1\}$, which corresponds to the cycle graph $C_{n}$ with vertices $V(C_{n})=\{x_{1},\dots,x_{n}\}$ and edges $E(C_{n})=\{\{x_{1},x_{2}\},\dots,\{x_{n-1},x_{n}\},\{x_{n},x_{1}\}\}$.
Figure 10. The circulant $C_{10}(2,5)$
Theorem 6.2.
Let $G=C_{n}=C_{n}(1)$. Then $G$ is levelable if and only if $n\in\{2,3,4,5,7\}$.
Proof.
One can easily check that $C_{n}$ is well-covered for $n\in\{2,3,4,5,7\}$ and hence levelable. For $n=6$, the maximal independent sets of $C_{6}$ are $\{1,3,5\},\{2,4,6\},\{1,4\},\{2,5\}$, and $\{3,6\}$. Taking $F_{1}=\{1,3,5\}$, $F_{2}=\{2,4,6\}$, $F_{3}=\{1,4\}$, and $F_{4}=\{2,5\}$, we see that the hypotheses of Lemma 2.7 are satisfied and so $C_{6}$ is not levelable.
We now show that $G=C_{n}$ is not levelable for $n\geq 8$. If $n\geq 8$ is even, we define
$F_{1}$ $=\{x_{1},x_{3},\dots,x_{n-1}\},~~F_{2}=\{x_{2},x_{4},\dots,x_{n}\},$
$F_{3}$ $=\{x_{1},x_{3},\dots,x_{n-5},x_{n-2}\},~~\mbox{and}~~F_{4}=\{x_{2},x_{4},\dots,x_{n-4},x_{n-1}\},$
and if $n\geq 9$ is odd, we define
$F_{1}$ $=\{x_{1},x_{3},\dots,x_{n-2}\},~~F_{2}=\{x_{2},x_{4},\dots,x_{n-1}\},$
$F_{3}$ $=\{x_{1},x_{3},\dots,x_{n-8},x_{n-5},x_{n-2}\},~~\mbox{and}~~F_{4}=\{x_{2},x_{4},\dots,x_{n-7},x_{n-4},x_{n-1}\}.$
In either case, one can check that the hypotheses of Lemma 2.7 are satisfied, and so it follows that $G$ is not levelable. ∎
A cubic circulant graph is a circulant graph such that every vertex has degree 3. We seek to classify those cubic circulants that are levelable. We begin with a structural result for cubic circulants due to Davis and Domke.
Theorem 6.3 ([7]).
Let $1\leq a<n$ and $t={\rm gcd}(2n,a)$.
  1. $(a)$
    If $\frac{2n}{t}$ is even, then $C_{2n}(a,n)$ is isomorphic to $t$ copies of $C_{\frac{2n}{t}}(1,\frac{n}{t})$.
  2. $(b)$
    If $\frac{2n}{t}$ is odd, then $C_{2n}(a,n)$ is isomorphic to $\frac{t}{2}$ copies of $C_{\frac{4n}{t}}(2,\frac{2n}{t})$.
By Lemma 2.1, it now suffices to determine for which $n\geq 2$ the graphs $C_{2n}(1,n)$ and $C_{2n}(2,n)$ are levelable. (Note that we may assume $n$ is odd for the latter.) Explicitly, these graphs have the vertex sets $V(C_{2n}(1,n))=\{x_{1},\dots,x_{2n}\}$ and $V(C_{2n}(2,n))=\{x_{1},\dots,x_{n},y_{1}\dots,y_{n}\}$ and edge sets
$E(C_{2n}(1,n))=\{\{x_{i},x_{i+1}\}\mid 1\leq i\leq 2n-1\}\cup\{\{x_{1},x_{2n}\}\}\cup\{\{x_{i},x_{n+i}\}\mid 1\leq i\leq n\}$
and
$E(C_{2n}(2,n))=$~ $\{\{x_{i},x_{i+1}\}\mid 1\leq i\leq n-1\}\cup\{\{y_{i},y_{i+1}\}\mid 1\leq i\leq n-1\}$
$\cup\{\{x_{1},x_{n}\},\{y_{1},y_{n}\}\}\cup\{\{x_{i},y_{i}\}\mid 1\leq i\leq n\}$
respectively. Figure 11 shows the labeling of these graphs will use in our argument.
$x_{n-2}$$x_{n-1}$$x_{n}$$x_{n+1}$$x_{n+2}$$x_{n+3}$$x_{3}$$x_{2}$$x_{1}$$x_{2n}$$x_{2n-1}$$x_{2n-2}$$x_{i}$$x_{n+i}$$x_{n-2}$$x_{n-1}$$x_{n}$$x_{1}$$x_{2}$$x_{3}$$y_{3}$$y_{2}$$y_{1}$$y_{n}$$y_{n-1}$$y_{n-2}$$x_{i}$$y_{i}$
Figure 11. The graphs $C_{2n}(1,n)$ and $C_{2n}(2,n)$ (with $n$ odd).
Theorem 6.4.
Let $G=C_{2n}(1,n)$. Then $G$ is levelable if and only if $n\in\{2,3,4\}$.
Proof.
For $n\in\{2,3,4\}$, one can directly check that the graphs $C_{2n}(1,n)$ are well-covered and hence levelable (also see [2, Theorem 4.3]). So suppose $n\geq 5$. We will show that $G=C_{2n}(1,n)$ is not levelable. We consider two cases.
Case 1: If $n\geq 5$ is odd, define
$F_{1}$ $=\{x_{1},x_{3},\dots,x_{2n-1}\},$
$F_{2}$ $=\{x_{2},x_{4},\dots,x_{2n}\},$
$F_{3}$ $=\{x_{1},x_{3},\dots,x_{n-4},x_{n-1}\}\cup\{x_{n+2},x_{n+4},\dots,x_{2n-3}\},$
$F_{4}$ $=\{x_{2},x_{4},\dots,x_{n-3},x_{n}\}\cup\{x_{n+3},x_{n+5},\dots,x_{2n-2}\}.$
We claim that $F_{1},\dots,F_{4}$ are maximal independent sets of $G$ that satisfy the hypotheses of Lemma 2.7.
Since $F_{3}\cup F_{4}\subsetneq F_{1}\cup F_{2}$ and $F_{3}\cap F_{4}\neq\emptyset$, it only remains to check that $F_{1},\dots,F_{4}$ are maximal independent sets of $G$. We see that $F_{1},\dots,F_{4}$ are independent sets by construction. Since $\{x_{1},x_{2n}\}$ and $\{x_{i},x_{i+1}\}$ are edges of $G$ for each $1\leq i\leq 2n-1$, it is clear that $F_{1}$ and $F_{2}$ are maximal. Additionally, since $\{x_{n-1},x_{2n-1}\}$ (resp. $\{x_{n},x_{2n}\}$) is an edge of $G$, we conclude that $F_{3}$ (resp. $F_{4}$) is maximal. It follows from Lemma 2.7 that $G$ is not levelable.
Case 2: If $n\geq 6$ is even, define
$F_{1}$ $=\{x_{1},x_{3},\dots,x_{n-1},x_{n+2},x_{n+4},\dots,x_{2n-2}\},$
$F_{2}$ $=\{x_{2},x_{4},\dots,x_{n},x_{n+3},x_{n+5},\dots,x_{2n-1}\},$
$F_{3}$ $=\{x_{1},x_{3},\dots,x_{n-5},x_{n-1},x_{2n-3}\}\cup A,$
$F_{4}$ $=\{x_{2},x_{4},\dots,x_{n-4},x_{n},x_{2n-2}\}\cup A^{\prime},$
where $A=A^{\prime}=\emptyset$ if $n=6$ and
$A$ $=\{x_{n+2},x_{n+4},\dots,x_{2n-6}\},$
$A^{\prime}$ $=\{x_{n+3},x_{n+5},\dots,x_{2n-5}\}$
for all $n\geq 8$.
We proceed as above. Since $F_{3}\cup F_{4}\subsetneq F_{1}\cup F_{2}$ and $F_{3}\cap F_{4}\neq\emptyset$, it only remains to check that $F_{1},\dots,F_{4}$ are maximal independent sets of $G$. We see that $F_{1},\dots,F_{4}$ are independent sets by construction. Since $\{x_{1},x_{2n}\}$ and $\{x_{i},x_{i+1}\}$ are edges of $G$ for each $1\leq i\leq 2n-1$, it is clear that $F_{1}$ and $F_{2}$ are maximal. Additionally, since $\{x_{n-3},x_{2n-3}\}$ and $\{x_{n-1},x_{2n-1}\}$ (resp. $\{x_{n-2},x_{2n-2}\}$ and $\{x_{n},x_{2n}\}$) are edges of $G$, we conclude that $F_{3}$ (resp. $F_{4}$) is maximal. It follows from Lemma 2.7 that $G$ is not levelable. ∎
Theorem 6.5.
Let $G=C_{2n}(2,n)$ with $n$ odd. Then $G$ is levelable if and only $n\in\{3,5\}$.
Proof.
For $n\in\{3,5\}$, one can directly check that the graphs $C_{2n}(2,n)$ are well-covered (also see [2, Theorem 4.3]) and hence levelable. So suppose $n\geq 7$ is odd. Define
$F_{1}$ $=\{x_{1},x_{3}\dots,x_{n-2}\}\cup\{y_{2},y_{4}\dots,y_{n-1}\},$
$F_{2}$ $=\{x_{2},x_{4}\dots,x_{n-1}\}\cup\{y_{3},y_{5}\dots,y_{n}\},$
$F_{3}$ $=\{x_{1},x_{3},\dots,x_{n-6},x_{n-3}\}\cup\{y_{2},y_{4},\dots,y_{n-5},y_{n-1}\},$
$F_{4}$ $=\{x_{2},x_{4},\dots,x_{n-5},x_{n-2}\}\cup\{y_{3},y_{5},\dots,y_{n-4},y_{n}\}.$
Since $F_{3}\cup F_{4}\subsetneq F_{1}\cup F_{2}$ and $F_{3}\cap F_{4}\neq\emptyset$, it only remains to check that $F_{1},\dots,F_{4}$ are maximal independent sets of $G$. We see that $F_{1},\dots,F_{4}$ are independent sets by construction. Since $\{x_{1},x_{n}\}$, $\{y_{1},y_{n}\}$, $\{x_{i},x_{i+1}\}$, and $\{y_{i},y_{i+1}\}$ are edges of $G$ for each $1\leq i\leq n-1$, it is clear that $F_{1}$ and $F_{2}$ are maximal. Additionally, since $\{x_{n-1},y_{n-1}\}$ and $\{x_{n-3},y_{n-3}\}$ (resp. $\{x_{n},y_{n}\}$ and $\{x_{n-2},y_{n-2}\}$) are edges of $G$, we conclude that $F_{3}$ (resp. $F_{4}$) is maximal. It follows from Lemma 2.7 that $G$ is not levelable. ∎
7. Levelable graphs and a connection to level artinian rings
This last section describes the origins of levelable graphs and their connection to level artinian rings. Level rings, which were introduced by Stanley [24] almost 50 years ago, have inspired a significant amount of research in combinatorial commutative algebra. As we shall show, our new connection allows us to use the properties of levelable graphs to deduce consequences for the existence of level artinian rings of a certain form.
Let $A$ be a graded $\mathbb{K}$-algebra, i.e., $A=\mathbb{K}[x_{1},\ldots,x_{n}]/I$ for some homogeneous ideal $I$. We will define level rings only in the case that $A$ is artinian – the general definition of level rings is in terms of the canonical module of $A$. The ring $A$ is artinian if the Krull dimension of $A$ is 0. Being artinian is equivalent to the fact that there exists positive integers $a_{1},\ldots,a_{n}$ such that $x_{i}^{a_{i}}\in I$ for $i=1,\ldots,n$. If $A$ is a graded artinian ring, then there is an integer $e$ such that
$A=A_{0}\oplus A_{1}\oplus A_{2}\oplus\cdots\oplus A_{e}~~\mbox{with $A_{e}\neq 0$}$
where each $A_{i}$ is a $\mathbb{K}$-vector space of homogeneous elements of $A$ of degree $i$. Note that $A_{i}=0$ for all $i>e$. The socle of $A$ is the homogeneous ideal of $A$ defined by
${\rm soc}(A)=\{a\in A~|~a\overline{x}_{i}=0~~\mbox{for $i=1,\ldots,n$}\},$
i.e., ${\rm soc}(A)$ is the annihilator of the maximal ideal $\langle\overline{x}_{1},\ldots,\overline{x}_{n}\rangle$ in $A$. The socle-vector is $s(A)=(s_{0},s_{1},\ldots,s_{e})$ where $s_{i}=\dim_{\mathbb{K}}{\rm soc}(A)_{i}$. We come to the key definition:
Definition 7.1.
A graded artinian algebra $A$ is level if the socle-vector of $A$ has the form $s(A)=(0,0,\ldots,0,s_{e})$ for some integer $s_{e}>0$.
In 2010, Van Tuyl and Zanello [26] introduced levelable simplicial complexes; these are simplicial complexes that allow one to build a level graded artinian ring. We recall the main ideas of Stanley–Resiner theory – see [14] for more details. A simplicial complex on a set $V=\{x_{1},\ldots,x_{n}\}$ is a subset of the power set of $V$ with the property that if $F\in\Delta$ and $G\subseteq F$, then $G\in\Delta$. The elements of $\Delta$ are called faces, and the maximal faces of $\Delta$ under inclusion are the facets. If $F_{1},\ldots,F_{s}$ is a complete list of the facets of $\Delta$, then we write $\Delta=\langle F_{1},\ldots,F_{s}\rangle$. The Stanley–Reisner correspondence allows one to associate a simplicial complex $\Delta$ on $V$ with a squarefree monomial ideal in $R=\mathbb{K}[x_{1},\ldots,x_{n}]$ with $\mathbb{K}$ as field. Precisely,
$I_{\Delta}=\langle x_{i_{1}}\cdots x_{i_{s}}~:~\{x_{i_{1}},\ldots,x_{i_{s}}\}\not\in\Delta\rangle.$
The ideal $I_{\Delta}$ is the Stanley–Reisner ideal of $\Delta$, and $R/I_{\Delta}$ is the Stanley–Reisner ring.
Given any simplicial complex and integers $a_{1},\ldots,a_{n}\geq 2$, set
$A(\Delta,(a_{1},\ldots,a_{n}))=\frac{\mathbb{K}[x_{1},\ldots,x_{n}]}{I_{\Delta}+\langle x_{1}^{a_{1}},\ldots,x_{n}^{a_{n}}\rangle}.$
The ring $A(\Delta,(a_{1},\ldots,a_{n}))$ is an example of a graded artinian ring. We omit the case that $a_{i}=1$ for some $i$ because in this situation
$A(\Delta,(a_{1},\ldots,a_{n}))\cong A(\Delta^{\prime},(a_{1},\ldots,\hat{a}_{i},\ldots,a_{n}))$
where $\Delta^{\prime}$ is the induced simplicial complex on $V\setminus\{x_{i}\}$.
In general, whether or not $A(\Delta,(a_{1},\ldots,a_{n}))$ is a level ring will depend upon both $\Delta$ and $(a_{1},\ldots,a_{n})$. Van Tuyl and Zanello called $\Delta$ a levelable simplicial complex if there exists $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ with all $a_{i}\geq 2$ such that $A(\Delta,(a_{1},\ldots,a_{n}))$ is a level ring. As shown by Van Tuyl and Zanello, determining if $\Delta$ is levelable reduces to solving a particular system of equations constructed from the facets of $\Delta$.
Theorem 7.2 ([26, Theorem 6]).
Let $\Delta$ be a simplicial complex on $V=\{x_{1},\ldots,x_{n}\}$ with facets $\{F_{1},\ldots,F_{s}\}$. Let $F_{i}=\{x_{i,1},\ldots,x_{i,d_{i}}\}$ denote the $i$-th facet. Then $A(\Delta,(a_{1},\ldots,a_{n}))$ is level if and only if $(a_{1},\ldots,a_{n})$ is an integral solution with $a_{i}\geq 2$ to the system
$(x_{1,1}+x_{1,2}+\cdots+x_{1,d_{1}})-(x_{2,1}+x_{2,2}+\cdots+x_{2,d_{2}})$ $=$ $d_{1}-d_{2}$
$(x_{2,1}+x_{2,2}+\cdots+x_{2,d_{2}})-(x_{3,1}+x_{3,2}+\cdots+x_{3,d_{3}})$ $=$ $d_{2}-d_{3}$
$\vdots$
$(x_{s-1,1}+x_{s-1,2}+\cdots+x_{s-1,d_{s-1}})-(x_{s,1}+x_{s,2}+\cdots+x_{s,d_{s}})$ $=$ $d_{s-1}-d_{s}.$
Levelable graphs, as we have defined them in this paper, arise when considering when the independence complex of a graph is a levelable simplicial complex. The independence complex of a graph $G$ is the simplicial complex
${\rm Ind}(G)=\{W\subseteq V~:~\mbox{$W$ is an independent set of $V$}\},$
that is, the faces of ${\rm Ind}(G)$ are the independent sets of $G$. The facets of ${\rm Ind}(G)$ is the set ${\rm MaxInd}(G)$ that we defined in the introduction. The connection between levelable graphs and level graded artinian rings is then caputerd in the following result.
Theorem 7.3.
A graph $G$ is a levelable graph if and only if ${\rm Ind}(G)$ is a levelable simplicial complex.
Proof.
Let ${\rm MaxInd}(G)=\{W_{1},\ldots,W_{s}\}$ be the maximal independent sets of $G$, and let $W_{i}=\{x_{i,1},\ldots,x_{i,d_{i}}\}$ for $i=1,\ldots,s$. Note that $W_{1},\ldots,W_{s}$ are also the facets of ${\rm Ind}(G)$.
(“Only If”) Suppose $G$ is a levelable graph. So there is a weight function $(c_{1},\ldots,c_{n})$ with positive integers and independence weight $c$ such that
$c_{i,1}+\cdots+c_{i,d_{i}}=c~\mbox{for $i=1,\ldots,s$}.$
If we now consider the vector $(c_{1}+1,\ldots,c_{n}+1)$, each $W_{i}$ gives the equality:
$(c_{i,1}+1)+\cdots+(c_{i,d_{i}}+1)=c+d_{i}.$
But then for $i=1,\ldots,s-1$, we have
$((c_{i,1}+1)+\cdots+(c_{i,d_{i}}+1))-((c_{i+1,1}+1)+\ldots+(c_{i,d_{i+1}}+1))=d_{i}-d_{i_{1}}.$
By Theorem 7.2, this means that ${\rm Ind}(G)$ is a levelable simplicial complexes since $(c_{1}+1,\ldots,c_{n}+1)$ is solution to the system of equations with $c_{i}+1\geq 2$ for all $i$.
(“If”) Let $(a_{1},\ldots,a_{n})$ with all $a_{i}\geq 2$ be an integral solution to the equations of Theorem 7.2 constructed from the facets of ${\rm Ind}(G)$. By reversing the above argument, we can show $G$ is levelable with respect to the weight function $(a_{1}-1,\ldots,a_{n}-1)$. ∎
In light of the previous result, finding levelable graphs immediately allows us to construct a level graded artinian ring. The edge ideal of $G$ is the squarefree monomial ideal $I(G)=\langle x_{i}x_{j}:\{x_{i},x_{j}\}\in E(G)\rangle$. It is well-known (e.g., see [14]) that $I(G)$ is the Stanley–Reisner ideal of the simplicial complex ${\rm Ind}(G)$, that is, $I(G)=I_{{\rm Ind}(G)}$. By simply translating definitions and using the previous theorem, we deduce Theorem 1.2:
Theorem 7.4.
Let $G=(V,E)$ be a graph with edge ideal $I(G)$ in $R=\mathbb{K}[x_{1},\ldots,x_{n}]$. Then $G$ is a levelable graph with respect to the weight function $(c_{1},\ldots,c_{n})$ if and only if
$R/(I(G)+\langle x_{1}^{c_{1}+1},\ldots,x_{n}^{c_{n}+1}\rangle)$
is a level graded artinian ring.
As the previous result shows, constructing level graded artinian rings from an edge ideal of a graph $G$ reduces to verifying that $G$ is a levelable graph. We conclude this paper by using a result of Brown and Nowakowski [4] to show that level graded artinian rings of this form are quite rare.
Recall from the introduction that we can associate to any $G$ a vector space
[…]
for a fixed field $F$. Note that to construct ${\rm WCW}(G)$ we must first fix a field. If the field has characterisic zero, then $F$ contains a copy of $\mathbb{N}$. We then have the following fact:
Lemma 7.5.
Fix a field $F$ of characteristic zero. Let $G$ be a graph and view ${\rm WCW}(G)$ as a vector space over $F$. If $\dim{\rm WCW}(G)=0$, then $G$ is not levelable.
Proof.
If $G$ was levelable, then there is a weight vector $(c_{1},\ldots,c_{n})$ with $c_{i}\geq 0$ and $c_{i}\in\mathbb{N}\subseteq F$. But then $(c_{1},\ldots,c_{n})\in{\rm WCW}(G)$ is a non-zero vector, and thus $0=\dim{\rm WCW}(G)\geq 1$, a contradiction. ∎
Example 7.6.
The converse of the previous statement is false. As we have seen, $P_{5}$ is not a levelable graph. However $(1,1,0,-1,-1)\in{\rm WCW}(P_{5})$, so $\dim{\rm WCW}(P_{5})\neq 0$.
We recall some of the terminology related to random graphs. A random graph $G$ of order $n$ is a graph on $V=\{x_{1},\ldots,x_{n}\}$ where there is an edge between $x_{i}$ and $x_{j}$ of fixed probability $p\in(0,1)$ for all possible $1\leq i<j\leq n$. We say that almost all graphs $G$ have property $P$ if ${\rm Prob}(\mbox{$G$ has property $P$})\rightarrow 1$ as $|V|\rightarrow\infty$.
Theorem 7.7 ([4, Theorem 1; Corollary 2]).
Fix a field $F$. For almost all graphs $G$, $\dim{\rm WCW}(G))=0$. In particular, for almost all graphs $G$, $G$ is not well-covered.
We can now deduce the following results that are of interest to those working in combinatorial commutative algebra.
Corollary 7.8.
For almost all graphs $G$, the graded artinian ring
$R/(I(G)+\langle x_{1}^{a_{1}},\ldots,x_{n}^{a_{n}}\rangle)$
is not level for any choice of $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ with $a_{i}\geq 2$ for all $i$.
Proof.
By Theorem 7.7, almost all graphs $G$ have $\dim{\rm WCW}(G)=0$, which by Lemma 7.5 implies that almost all graphs are not levelable, and thus by Theorem 7.4, for almost all graphs, the ring in the statement of the corollary cannot be level. ∎
One class of rings that are of great importance in commutative algebra are Cohen–Macaulay rings. We say a ring $S$ is Cohen–Macaulay if $\dim S={\rm depth}S$. A graph $G$ is a Cohen–Macaulay graph if the ring $R/I(G)$ is a Cohen–Macaulay ring (for more details, see [14]). For $G$ to be Cohen–Macaulay, a necessary property is that $G$ is well-covered. Consequently, we can also deduce the following result from Theorem 7.7:
Corollary 7.9.
For almost all graphs $G$, $G$ is not Cohen–Macaulay.
Remark 7.10.
The previous result was also proved by Docthermann–Newman [8, Corollary 1.4] and Erman–Yang [9, Corollary 7.1], using approaches different than Brown–Nowakowski.
Acknowledgments
Some of our results first appeared in the senior undergraduate thesis of M. Chong under the supervision of A. Van Tuyl. We thank David Tankus for his comments, and in particular, for his improvement to Corollary 3.8 and bringing [20] to our attention. Part of this work was carried out while T. Hibi visited the other authors at McMaster University, and when T. Hibi and A. Van Tuyl visited the Fields Institute in Toronto as part of the “Thematic Program in Commutative Algebra and Applications”. We thank the hospitality of both institutions. A. Van Tuyl’s research is supported by NSERC Discovery Grant 2024-05299.